OpenStudy (anonymous):

consider the iterated triple integral a. sketch the region of integration b. express with order of integration dxdydz c. express with order of integration dydxdz

OpenStudy (anonymous):

o gahd @jim_thompson5910

OpenStudy (anonymous):

I always hated these type of problems

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{1-x}^{1}\int\limits_{x}^{1}f(x,y,z)dzdydx\]

OpenStudy (anonymous):

haha me too

OpenStudy (anonymous):

no function?

OpenStudy (anonymous):

o well first you need to graph each region

OpenStudy (anonymous):

i'll try to help you as much as i can but i realy didn't understand this when i did it... @dpaInc you remember this?

OpenStudy (anonymous):

first draw your regions on the xy,xz,yx planes

OpenStudy (turingtest):

you only need one region here I think

OpenStudy (anonymous):